40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2676     Entry time: Mon Mar 8 12:11:15 2021
Author: aaron 
Type: DailyProgress 
Category: Laser 
Subject: busby, sr554 to cryo lab 

enter Mon Mar 8 12:11:07 2021

Making the Allan covariance measurement today.

Picked up the busby low noise box and SR554 from gautam at 40m, which I'll use to measure the current noise of our laser drivers.

I returned the fiber components to the previous state after measurement (PSOMA S beam incident on cavity, N beam to the other fiber launch, both beams pass through their respective EOM).

exit Mon Mar 8 19:22:50 2021

ELOG V3.1.3-